OpenStudy (anonymous):

evaluate using product (x^2-1)^3*(x^2+2)

OpenStudy (anonymous):

prouct rule?

OpenStudy (anonymous):

it sasys i can use product, quotient or chain???

OpenStudy (bahrom7893):

\[y = (x^2 - 1)^3 * (x^2+2)\]

OpenStudy (anonymous):

you got this bahrom! lol

OpenStudy (bahrom7893):

\[y' = (x^2-1)^3 \] \[y' = (x^2-1)^3*(2x) + (x^2+2)*3(x^2-1)^2*(2x)\]

OpenStudy (bahrom7893):

sorry ignore the first equation, i forgot to erase.

OpenStudy (bahrom7893):

\[y' = (x^2-1)^2 [2x(x^2-1) + 6x(x^2+2)]\]

OpenStudy (bahrom7893):

\[y' = (x^2-1)^2 2x[(x^2-1) + 3(x^2+2)]\]

OpenStudy (bahrom7893):

\[y' = 2x(x^2-1)^2 [(x^2-1) + 3(x^2+2)]\]